OpenStudy (anonymous):

Find the equation of the line parallel to 4x+3y-15=0 that has an x intercept of -3 in standard form.

OpenStudy (anonymous):

First, change the 4x+3y-15=0 to slope-intercept form (y = mx+b) What do you get?

OpenStudy (anonymous):

-4/3x+5=y

OpenStudy (anonymous):

Good, now do you know what parallel lines have in common (slope)

OpenStudy (anonymous):

@zaynahf Their slope.

OpenStudy (anonymous):

Goodyyyy.. if the slope of that is -4/3, and its parallel, then the slope of the new one will also be -4/3. And if the x-int is (-3,0), you can use that in point-slope form. Do you know point-slope form?

OpenStudy (anonymous):

Oh I got the answer, thanks!